# Dialogue Systems for Emotional Support via Value Reinforcement

### v1

@@ 6.3 Baselines
We evaluate our approach against three categories of baseline models:
  • Prompt-Based Models: GPT-4o-mini and Llama-3-8B-Instruct.
  • Fine-Tuned Models: Variants of Llama-3-8B-Instruct trained on existing emotional support datasets, including ESConv (Liu et al., 2021), ExTES (Zheng et al., 2024), and Psych8k (Liu et al., 2023). These datasets are based on English-language conversations.
  • Emotion-Reinforced Models: To verify the effectiveness of value reinforcement, we train the reference generator and supporter model to prioritize positive emotion reinforcement instead of values (Appendix G).
@@
Fine-tuned Llama-ESConv 4.35 4.43 4.06 2.65 3.88 4.82 4.97 3.79 4.25 4.22       1.87 0.37       0.19
@@
6.4 Evaluation Results
6.4.1 Effectiveness of Value Targeting and Reference Responses
As shown in Table 2, leveraging both target value prediction and reference responses significantly improved performance across all ES-Skills metrics while reducing ES-Intensity. This approach notably enhanced key ES-Skills, including Suggesting, Expression, and Informativeness, resulting in more balanced emotional support. Similarly, value reinforcement was substantially improved when both target values and reference responses were utilized. These findings emphasize the effectiveness of targeting specific values at each turn and using reference responses that leverage real-world knowledge from Reddit. Since our fine-tuned models are trained on GPT-simulated data, we use the simulation data that incorporates both target values and reference responses in subsequent experiments.
@@ 6.4.2 Performance Comparison with Baselines
The performance comparisons between our models and the baselines are presented in Tables 3. For our DPO model and the emotion-reinforced model (DPO), we select configurations based on their optimal performance ($h=3$, $\gamma=1$, $T_{\text{diff}}=2$ and $h=3$, $\gamma=1$, $T_{\text{diff}}=0.5$, respectively). For results with more DPO hyperparameters, refer to Table 20 in the Appendix.
@@
Our DPO model (row 9) outperformed both prompt-based and fine-tuned baselines across most metrics, particularly in emotional support metrics such as Suggestions, Experience, and Informativeness. These improvements highlight the characteristics of our reference responses, which focused on sharing relevant personal experiences, providing practical solutions, and fostering self-confidence. These are key components of effective online emotional support. The models also showed significant gains in conversational capabilities, particularly in the Expression and Humanness metrics, leading more natural and dynamic interactions.
Notably, the variant of our method that focuses on reinforcing positive emotions rather than values (rows 6–7) also consistently outperformed other baselines. This suggests that one of our key ideas—leveraging crowd knowledge from Reddit—is still effective when the supporter model is designed to promote positive emotions in seekers. Yet, our value reinforcement approach achieved higher scores across most emotional support skill metrics at comparable training stages, highlighting the greater effectiveness of reinforcing values over positive emotions in enhancing emotional support.
@@
Our DPO model (row 9) outperformed most baselines, demonstrating that our supporter model reduces the intensity of seekers’ negative emotions more effectively than other methods. Interestingly, our models (rows 7–8) also resulted in lower intensity than the emotion reinforcement models at comparable training stages (rows 6–7). This suggests the potential of alleviating a seeker’s negative emotions indirectly by redirecting their focus to values, presenting this indirect strategy as a promising research direction for future work.
Llama-Psych8k showed significantly lower ES-Intensity than our model. Analysis revealed it generated much longer responses (73 words per turn) compared to our model and other baselines (20–25 words). Since the ES-Intensity model was validated on ESConv, caution is warranted when interpreting the scores of dialogues with substantially different distributions. Moreover, in practice, its lengthy responses and lower Humanness scores may feel overwhelming, discouraging seeker engagement.
@@
Despite the strong performance of our DPO model compared to most baselines, its results were comparable to those of the emotion-reinforcing DPO and certain fine-tuned models. An analysis of 40 dialogues with low ES-Value scores revealed two key areas for improvement: enhancing the ability to identify seekers’ unique strengths and accomplishments, and improving the capacity to address their emotional states and concerns more deeply. These findings underscore the need to improve the model’s engagement with seekers’ individual circumstances, which is expected to enhance its value reinforcement performance. Detailed experimental methods and results are provided in Appendix J.
@@ 6.4.3 Expert Evaluation
Another strength is the model’s capacity to emphasize positive aspects of the seeker’s situation. For example, responses like “Your initiative to seek meaningful experiences reflects your dedication to making a difference, and that determination will surely lead you closer to your goals.” help seekers recognize their strengths and positive attributes. This approach enhances the seeker’s self-awareness and supports the reinforcement of their values in a meaningful and constructive way.
@@ 7 Conclusion
In this paper, we introduce an emotional support framework based on value reinforcement, designed to promote long-term emotional well-being. The framework incorporates a target value detector and a reference generator to improve the supporter model’s ability to generate value-aligned and effective support responses. Evaluations demonstrate that our framework surpasses baseline models in both emotional support quality and value reinforcement. Expert therapist evaluations further highlight the model’s strengths in validating seekers’ challenges and emphasizing positive aspects of their situations, which are key elements of effective emotional support. These results underscore the potential of value reinforcement to enhance supportive interactions and provide a foundation for developing more effective emotional support systems.
@@
Sharing experiences related to those of the seeker- is a key strategy in emotional support for fostering intimacy and has been a key evaluation criterion in prior emotional support systems (Zhang et al., 2023). However, some users might feel uncomfortable when dialogue systems present these experiences as personal. We found that removing self-disclosure strategy from the model impacts the quality of emotional support (Appendix I), highlighting the need for further research into more sophisticated approaches to experience sharing, which we leave as a direction for future work.
@@ Appendix A Experiment Details for Identifying the Effect of Values in Emotional Support
Since we are interested in cases where values are expressed positively, we only count the values when the seeker’s utterances are positive. To determine whether the seeker’s utterances are positive, we use the EmoLlama-Chat-7B proposed by Liu et al. (2024). To extract the values in the seeker’s utterances, we employ the model that achieved the best performance in the SemEval2023 Task 4: Identification of Human Values behind Arguments (Schroter et al., 2023; Kiesel et al., 2023).
Appendix B Dataset Settings in Reddit
@@ E.1 Persona Generation
Our persona generation process resulted in 2,036 unique personas, each defined by problem category, human values, situations, emotion types, and demographic information. These personas are divided into three datasets: a training set containing 1,796 personas, and development and test sets with 120 personas each. The training and development sets are used to construct SFT and DPO datasets for the supporter model through simulation, while the test set is reserved for comparative performance evaluation across models. Examples of the generated personas are provided in Table 27.
E.2 Evaluation of Seeker Simulator Performance
@@ Results.
E.3 Prompts for Seeker Simulator
@@ F.3 ES-Value
Appendix G Training Details for Emotion-Reinforced Models
@@ G.2 Supporter Model
Appendix H Validation of Automated Evaluation Models
@@
Appendix I Details of Experiments on Self-Disclosure
@@
Appendix J Detailed Analysis of Value Reinforcement Performance
@@ 4. Reference response: {Reference response}
Fine-tuned Llama-ESConv - - - 4.35 4.43 4.06 2.65 3.88 4.82 4.97 3.79 4.25 4.22       1.87
@@
  Fine-tuned Llama-ESConv - - - 0.37       0.21 0.35       0.17 0.37       0.19

### v2

@@ 6.3 Baselines
  • Prompt-Based: GPT-4o-mini and Llama-3-8B-Instruct.
  • ES Datasets: Variants of Llama-3-8B-Instruct trained on emotional support datasets, including Reddit (Section 4), ESConv (Liu et al., 2021), ExTES (Zheng et al., 2024), and Psych8k (Liu et al., 2023).
  • ES Methods: Recent methods contributing to the development of supporter models, including Ask-an-Expert (Zhang et al., 2023), ESCoT (Zhang et al., 2024), and PPDPP (Deng et al., 2024). Details of these methods are provided in Appendix G.
  • Emotion-Reinforced: To verify the effectiveness of value reinforcement, we train the reference generator and supporter model to reinforce positive emotions instead of values (Appendix H).
6.4 Evaluation Results
6.4.1 Effectiveness of Value Targeting and Reference Responses
As shown in Table 2, leveraging both target values and reference responses significantly improved performance across all ES-Skills metrics while reducing ES-Intensity. This approach notably enhanced key ES-Skills, including Suggesting, Expression, and Informativeness. Similarly, value reinforcement was substantially improved when both target values and reference responses were utilized. These findings emphasize the effectiveness of targeting specific values at each turn and using reference responses that leverage real-world knowledge from Reddit. Since our fine-tuned models are trained on GPT-simulated data, we use the simulation data that incorporates both target values and reference responses in subsequent experiments.
@@ 6.4.1 Effectiveness of Value Targeting and Reference Responses
ES Datasets Llama-Reddit 3.38 3.74 3.21 2.59 2.99 3.94 4.35 3.37 3.81 3.40       1.97 0.29       0.09
Llama-ESConv 4.35 4.43 4.06 2.65 3.88 4.82 4.97 3.79 4.25 4.22       1.87 0.37       0.19
@@ 6.4.2 Performance Comparison with Baselines
The performance comparisons between our models and the baselines are presented in Table 3. For our DPO and emotion-reinforced DPO models, we select optimal configurations ($h=3$, $\gamma=1$, $T_{\text{diff}}=2$ and $h=3$, $\gamma=1$, $T_{\text{diff}}=0.5$, respectively). For results with more DPO hyperparameters, refer to Table 24 in the Appendix.
@@
Our DPO model outperformed the baselines across most metrics, particularly in emotional support metrics such as Suggestions, Experience, and Informativeness. These improvements reflect the characteristics of our reference responses, which emphasize sharing relevant experiences and offering practical solutions—key elements of effective online emotional support. The models also demonstrated significant gains in conversational capabilities, especially in Expression and Humanness, resulting in more natural and engaging interactions.
Notably, the variant of our method that focuses on reinforcing positive emotions rather than values (Emotion-Reinforced) also consistently outperformed other baselines. This suggests that one of our key ideas—leveraging crowd knowledge from Reddit—is still effective when the supporter model is designed to promote positive emotions in seekers. Yet, our value reinforcement approach achieved higher scores across most emotional support skill metrics at comparable training stages, highlighting the effectiveness of reinforcing values in enhancing emotional support.
@@
Our DPO model outperformed most baselines, demonstrating that our supporter model reduces the intensity of seekers’ negative emotions more effectively than other methods. Notably, our models achieved lower intensity levels than the emotion reinforcement models at comparable training stages. This suggests that redirecting a seeker’s focus to values can indirectly alleviate negative emotions, highlighting a promising direction for future research.
Llama-Psych8k showed significantly lower ES-Intensity than our model. Analysis revealed it generated much longer responses (73 words per turn) compared to other models (20–25 words). Since the ES-Intensity model was validated on ESConv, caution is warranted when interpreting the scores of dialogues with substantially different distributions. Moreover, in practice, its lengthy responses and lower Humanness scores may feel overwhelming, discouraging seeker engagement.
@@
Despite the strong performance of our DPO model compared to most baselines, its results were comparable to those of the emotion-reinforcing DPO and certain fine-tuned models. An analysis of 40 dialogues with low ES-Value scores revealed two key areas for improvement: enhancing the ability to identify seekers’ unique strengths and accomplishments, and improving the capacity to address their emotional states and concerns more deeply. These findings underscore the need to improve the model’s engagement with seekers’ individual circumstances, which is expected to enhance its value reinforcement performance. Detailed experimental methods and results are provided in Appendix M.
Categories Models Valid Turns
1 2 3
Prompt- Based GPT 0.681 0.728 0.740
Llama 0.688 0.728 0.751
Ours DPO 0.720 0.760 0.779
Table 4: Average success rates of target value reinforcement within 1, 2, and 3 valid turns—the number of future turns within which the target values remained relevant.
6.4.3 Success of Target Value Reinforcement
As our approach focuses on setting target values and reinforcing them, we evaluated its effectiveness in reinforcing the target values. To this end, we compared our models with other prompt-based models that also take target values and reference responses as inputs.
Table 4 presents the average success rates across models, showing that our model consistently outperforms others in all cases for the next 1–3 turns. Since all models use the same inputs—target values and a reference response—these differences underscore the effectiveness of our supporter model’s training methods in reinforcing target values. Further details of this analysis are provided in Appendix I.
6.4.4 Generalization beyond Reddit Data
To evaluate the generalizability of our target value detection and reference generation methods beyond Reddit, we explored training these two models on the Cactus dataset (Lee et al., 2024)—counseling conversations based on Cognitive Behavioral Theory (CBT)—instead of Reddit. Detailed procedures are provided in Appendix J.
For ES-Value (shown in Table 26 for clarity), DPO (Cactus) outperformed all baselines, demonstrating the supporter model’s ability to effectively learn and apply value signals from Cactus. This result confirms the generalizability of our target value detection and reference response generation approach.
Similarly, for ES-Skills (Table 3), DPO (Cactus) performed comparably to or better than the baselines, particularly excelling in emotional support metrics. However, it scored significantly lower on the Experience metric compared to the original DPO model. Since Reddit contains numerous shared personal narratives, the reference response generator appears to benefit more from Reddit than from Cactus, supporting our decision to use Reddit.
@@ 6.4.5 Expert Evaluation
Another strength is our model’s capacity to emphasize positive aspects of the seeker’s situation, reflecting positive values and related goal. For example, responses like “Your initiative to seek meaningful experiences reflects your dedication to making a difference, and that determination will surely lead you closer to your goals.” help seekers recognize their strengths and positive attributes. This behavior was contrasted with GPT-4o’s responses, particularly when the seeker persisted in a negative mood. GPT-4o tended to focus heavily on expressing empathy and lingered in the negative mood. This overemphasis on empathy is likely a result of the human preference alignment process. Our method overcomes this tendency by enhancing the seeker’s self-awareness and supporting the reinforcement of their values in a constructive manner.
@@ 7 Conclusion
In this paper, we introduce the first emotional support framework based on value reinforcement, as emphasized in modern psychotherapy. The framework incorporates a target value detector and a reference generator to improve the supporter model’s ability to generate value-aligned and effective support responses. Evaluations demonstrate that our framework surpasses baseline models in both emotional support quality and value reinforcement. Expert therapist evaluations further highlight the model’s strengths in validating seekers’ challenges and emphasizing positive aspects of their situations, which are key elements of effective emotional support. These results underscore the potential of value reinforcement to enhance supportive interactions and provide a foundation for developing more effective emotional support systems.
@@
Sharing experiences related to those of the seeker is a key strategy in emotional support for fostering intimacy and has been a key evaluation criterion in prior emotional support systems (Zhang et al., 2023). However, some users might feel uncomfortable when dialogue systems present these experiences as personal. We found that removing self-disclosure strategy from the model impacts the quality of emotional support (Appendix L), highlighting the need for further research into more sophisticated approaches to experience sharing, which we leave as a direction for future work.
@@ Appendix A Experiment Details for Identifying the Effect of Values in Emotional Support
In this experiment, we compare positive value expressions in the seeker’s final four turns between high effectiveness and low effectiveness groups. This analysis window is chosen based on three key considerations: (1) To control for variability in dialogue length within ESConv, we analyze a fixed number of turns. (2) To capture changes resulting from emotional support conversations, we focus on the latter part of the dialogue. (3) Given that the minimum number of turns in ESConv is eight, we set the analysis window at half this value.
Since we are interested in cases where values are expressed positively, we count the values only when the seeker’s utterances are positive. To determine whether an utterance is positive, we use the EmoLlama-Chat-7B proposed by Liu et al. (2024). To extract the values in the seeker’s utterances, we employ the model that achieved the best performance in the SemEval2023 Task 4: Identification of Human Values behind Arguments (Schroter et al., 2023; Kiesel et al., 2023).
Appendix B Constructing the Emotional Support Dataset from Reddit
B.1 Models for Sentiment Strength and Values Labeling
Automated classification models are essential for labeling Reddit data at scale. We prioritize models with optimal performance, as classification accuracy directly influences the quality of subsequent fine-tuning.
Sentiment Strength.
We employ EmoLlama-Chat-7B (Liu et al., 2024), which demonstrates superior performance in the V-reg task, estimating emotional valence on a continuous scale from 0 (most negative) to 1 (most positive). The comparative performance of various models, including both fine-tuned models and zero/few-shot methods, is presented in Table 5.
Human Values.
We adopt the top-performing model (Schroter et al., 2023) from SemEval 2023 Task 4: Identification of Human Values behind Arguments (Kiesel et al., 2023), a shared task highly relevant to our objective based on the same value categorization. Considering recent advancements in LLMs, we also compare its performance against GPT-4o-mini by experimenting with two strategies: (1) predicting the binary presence of each value for a given sentence, and (2) assigning a probability score (between 0 and 1) to indicate the level of support for each value. As shown in Table 6, evaluation on the shared task’s test dataset demonstrates that the model from the shared task consistently outperforms GPT-4o-mini, supporting our decision to adopt it as our classification model. For reference, the F1-score of random classification is 0.128.
Category Model Corr.
Fine-tuned BERT-base 0.840
RoBERTa-base 0.845
SentiBERT 0.835
Zero-shot/ Few-shot Falcon 0.135
Vicuna 0.298
LLaMA2-7B-chat 0.094
LLaMA2-13B-chat 0.312
ChatGPT 0.637
ChatGPT-FS 0.739
GPT-4 0.811
GPT-4-FS 0.825
Instruction-tuned EmoLlama-Chat-7B 0.876
Table 5: Performance comparison of models on the V-reg task, measured using Spearman correlation coefficients. Results are from Liu et al. (2024)
Models F1-Score
Model from Shared Task 0.57
GPT (Binary) 0.38
GPT (Prob.) 0.40
Table 6: Performance comparison of the best-performing shared task model and GPT-4o-mini in a multi-label classification task across 20 values.
B.2 Dataset Settings in Reddit
@@ E.1 Persona Generation
Our persona generation process resulted in 2,036 unique personas, each defined by problem category, situations, emotion types, and demographic information. These personas are divided into three datasets: a training set containing 1,796 personas, and development and test sets with 120 personas each. The training and development sets are used to construct SFT and DPO datasets for the supporter model through simulation, while the test set is reserved for comparative performance evaluation across models. Examples of the generated personas are provided in Table 32.
E.2 Evaluation of Seeker Simulator Performance on ESConv
@@ Results.
E.3 Human Evaluation of Seeker Simulator
We conducted a human evaluation to assess the naturalness of responses generated by our seeker simulator compared to real human seekers. The evaluation utilized two psychotherapy datasets: (1) ESConv (Liu et al., 2021), a crowdsourced emotional support dialogue dataset, and (2) AnnoMI (Wu et al., 2022, 2023), which comprises real counseling conversations from YouTube and Vimeo videos.
The evaluation sample included 200 dialogues (140 from ESConv and 60 from AnnoMI) truncated to various lengths. For each dialogue history, we compared the seeker simulator’s generated responses with the original seeker responses in a pairwise manner.
The evaluation was conducted by 16 evaluators, comprising undergraduate and graduate students from diverse academic backgrounds, including psychology, education, and computer science. To ensure an unbiased assessment, they had no prior exposure to our system. The evaluators were asked, “Which response is more natural for the seeker?” and instructed to choose between the two responses or select “Tie.” While the evaluators were not psychotherapy experts, they could reliably assess response naturalness based on their own experiences with help-seeking situations, as emotional distress is a universal human experience.
The results, presented in Figure 4, show that the seeker simulator’s responses were rated as more natural than the original seeker responses in 66.4% and 33.3% of cases across the two datasets. Including ties, these percentages increased to 80.7% and 75.0%. Some evaluators noted that the seeker simulator effectively conveyed negative emotions and demonstrated strong situational engagement. These findings suggest that the seeker simulator achieves a level of naturalness comparable to that of real seekers.
E.4 Prompts for Seeker Simulator
@@ F.3 ES-Value
Appendix G Descriptions of Emotional Support Methods Selected as Baselines
We select baselines from methods that have contributed to the development of supporter models, ensuring diverse characteristics. The explanations for each method are as follows:
  • Ask-an-Expert (Zhang et al., 2023): This approach involves consulting an expert-role LLM at every turn of the conversation to obtain advice on the seeker’s emotional status, its cause, and potential solutions, which are then leveraged to generate the supporter’s response.
  • ESCoT (Zhang et al., 2024): This method follows a chain-of-thought process, considering the seeker’s emotional state, emotion stimulus (the specific trigger of the emotion), the seeker’s personal interpretation of the stimulus, and the reasoning behind the selected support strategy before generating a supporter response.
  • PPDPP (Deng et al., 2024): This approach involves training a policy planner to select the optimal support strategy through two steps: (1) fine-tuning on the ESConv dataset and (2) simulating and evaluating diverse conversations using three LLMs (supporter LLM, seeker LLM, and reward LLM), followed by reinforcement learning based on the rewards.
Appendix H Training Details for Emotion-Reinforced Models
@@ H.2 Supporter Model
Appendix I Evaluating the Success of Target Values Reinforcement
We engaged each model in conversations with a seeker simulator and analyzed the frequency of target values appearing in the seeker’s subsequent responses. The success rate was assessed based on valid turns—the number of future turns within which the target values remained relevant. For example, if the valid turn threshold was three, reinforcement was considered successful if the seeker’s response included the target values within the next three turns. To control for variations in dialogue length, we considered only cases where at least one positive seeker response occurred within the valid turns. This approach mitigated the influence of longer conversations, where negative responses might become more frequent and lower the success rate.
Appendix J Training Details for Generalization Capability Evaluation
To evaluate the generalization capability of our method, we conducted experiments using the Cactus dataset (Lee et al., 2024) instead of Reddit. Cactus is a counseling dataset based on Cognitive Behavioral Therapy (CBT), a therapeutic approach that helps individuals identify and modify negative thought patterns and behaviors. The conversations were generated using GPT-4o and validated through human evaluation to ensure their suitability as psychological counseling dialogues.
We utilized Cactus to train both the target value detector and the reference generator. Subsequently, we employed these models in a simulation to create SFT and DPO datasets, which were then used to train the supporter model. The details for each stage are outlined below.
J.1 Target Value Detector & Reference Generator
The Cactus dataset contains a total of 31,577 counseling dialogues. To consider resource efficiency, we selected 4,057 dialogues by excluding cases where the seeker’s intake form—including personal information, issues, history, and other details—was duplicated across conversations. As described in Section 4, we then used models developed by Liu et al. (2024) and Schroter et al. (2023) to label each seeker’s turn with sentiment strength and expressed values. These results were subsequently used to prepare the dataset for training the target value detector and reference generator. The final dataset consists of 11,000 instances for training and 1,631 instances each for validation and testing.
Following the approach outlined in Section 5, we trained the target value detector and reference generator separately. Unlike Reddit, where a single comment can receive multiple replies—leading to a branching conversation structure—the Cactus dataset follows a unidirectional conversational flow. Due to this structure, the DPO process was omitted during reference generator training.
Stage Supporter Train Dev
SFT GPT-4o-mini 38,543 2,640
DPO GPT-4o-mini 1,270 256
Table 16: Dataset sizes for supporter model training, generated through simulation using Cactus-based target value detector and reference generator. The ‘Supporter’ column refers to the supporter model used in the simulation.
J.2 Supporter Model
We trained the supporter model in two stages, SFT and DPO, following the approach introduced in Section 5.3. The training dataset was generated through interactions between GPT-4o-mini-based supporter and seeker simulators. The key difference from Section 5.3 is that the target value detector and reference generator used for simulation were trained on the Cactus dataset instead of Reddit. Additionally, when constructing the dataset for DPO, we adopted the hyperparameter settings ($h=3$, $\gamma=1$, $T_{\text{diff}}=2$) that achieved the best performance on the Reddit dataset. Detailed information on dataset size is provided in Table 16.
@@ Appendix I Evaluating the Success of Target Values Reinforcement
Appendix K Validation of Automated Evaluation Models
@@
Appendix L Details of Experiments on Self-Disclosure
@@
Appendix M Detailed Analysis of Value Reinforcement Performance
@@ 4. Reference response: {Reference response}
ES Datasets Llama-Reddit - - - 3.38 3.74 3.21 2.59 2.99 3.94 4.35 3.37 3.81 3.40       1.97
Llama-ESConv - - - 4.35 4.43 4.06 2.65 3.88 4.82 4.97 3.79 4.25 4.22       1.87
@@
   ES Datasets Llama-Reddit - - - 0.30       0.10 0.28       0.07 0.29       0.09
Llama-ESConv - - - 0.37       0.21 0.35       0.17 0.37       0.19
@@
[…]
Seeker Supporter
Prompt- Based GPT 0.46       0.28
Llama 0.48       0.42
ES Datasets Llama-Reddit 0.24       0.04
Llama-ESConv 0.30       0.11
Llama-ExTES 0.44       0.37
Llama-Psych8k 0.46       0.45
ES Methods Ask-an-Expert 0.28       0.08
ESCoT 0.24       0.03
PPDPP 0.40       0.20
Ours DPO (Cactus)        -         -
Table 26: Comparison of models based on ES-Value—the win ratio of each model against DPO (Cactus). A value below 0.5 indicates that DPO (Cactus) outperformed the baseline models, while a value above 0.5 suggests that the baselines performed better. Statistically significant differences compared to DPO (Cactus) are marked with * ($p$-value $<0.05$), and differences with $p$-value $<0.1$ are marked with $\dagger$, as determined by the Mann-Whitney U test.